OpenStudy (anonymous):

185/100 in lowest terms and decimal

OpenStudy (anonymous):

1.85

OpenStudy (anonymous):

what is 185/100 in lowest terms too please

OpenStudy (anonymous):

37/20

OpenStudy (anonymous):

185/100 = 37/20 = 1.85